Sales of tablet computers at Ted Glickman's electronics store in Washington,D,C., over the past 10 weeeks are shown in the table below:

Week Damand

1 20

2 21

3 28

4 37

5 25

6 29

7 36

8 22

9 25

10 28

a)Forecast demand for each week, including week 10, using exponential smoothing with a=0.5 (initial forecast=20)

b)Compute the MAD

c)Compute the tracking signal
Answer #1
Smoothing Factor 0.5
Week 1 Week 2 Week 3 Week 4 Week 5 Week 6 Week 7 Week 8 Week 9 Week 10 Total
Demand(D) 20 21 28 37 25 29 36 22 25 28
Forecast(F) 20 20 20.5 24.25 30.625 27.8125 28.40625 32.20313 27.10156 26.05078
D - F 0 1 7.5 12.75 -5.625 1.1875 7.59375 -10.2031 -2.10156 1.949219 14.05078
|D - F| 0 1 7.5 12.75 5.625 1.1875 7.59375 10.20313 2.101563 1.949219 49.91016
MAD 4.991016
Tracking Signal 2.815215

Formulae :

1. Forecast = (((1 - smoothing factor) * Most recent period forecast) + (Smoothing factor * most recent periods' demand)

2. MAD =

where N = Number of period ( Here its 10)

3. Tracking Signal =

Example:

Week 1, Forecast = 20 (given in question)

Week 2, Forecast = (((1 - 0.5) * 20) + (0.5 * 20)) = 20

Week 3, Forecast = ((( 1 - 0.5) * 20) + (0.5 * 21)) = 20.5

........

MAD = (49.91016/ 10) = 4.991016

Tracking Signal = (14.05078 / 4.991016) = 2.815215